OpenStudy (camper4834):

At what point does the curve have the maximum curvature? y=7ln(x) I know how to find the curvature but after finding the formula for the curvature I have to use that to get the critical points and the maximum... it is all very tedious. I was wondering if there was a better way.

OpenStudy (camper4834):

@dan815

OpenStudy (camper4834):

@Hero

OpenStudy (anonymous):

what is you're way you started love

OpenStudy (camper4834):

\[k=\frac{\frac{ -7 }{ x^2 }}{\sqrt{1+\frac{49}{x^2}}^3}\]

OpenStudy (camper4834):

now to find the greatest curvature i need to take the derivative of that and set it equal to zero. then solve for x. do you see why this is a problem?

OpenStudy (anonymous):

mhm hold on love i am thinkingi might be in corrcet about what i am ghoing t otype bhut i going to ask my friend to help wme explan it to you @jabez177

OpenStudy (anonymous):

@jabez177 can you help me explan to this user?

OpenStudy (anonymous):

never mind yes i do see the probleom

jabez177 (jabez177):

Hello! Too many people tagging me.

OpenStudy (camper4834):

hi jabez. in this problem can solve it but not without a calculator. is there a method you can show me where it is possible to solve by hand?

OpenStudy (anonymous):

First of all, k(t) = |f ''(x)| / (1 + (f '(x))^2)^(3/2) ......= |-1/x^2| / (1 + (1/x)^2)^(3/2) ......= (1/x^2) / [(1/x^2) (x^2 + 1)]^(3/2) ......= (1/x^2) / [(1/x^3) (x^2 + 1)^(3/2)] ......= x(x^2 + 1)^(-3/2). Now, we can find the maximal curvature. k'(t) = 1(x^2 + 1)^(-3/2) + x * -3x(x^2 + 1)^(-5/2) ......= (x^2 + 1)^(-5/2) [(x^2 + 1) - 3x^2], by factoring ......= (x^2 + 1)^(-5/2) (1 - 2x^2). Setting k' = 0 for critical points yields 1 - 2x^2 = 0 ==> x = ±1/√2. Since k' > 0 for x in (-1/√2, 1/√2), and k' < 0 for x < -1/√2 or x > 1/√2, we conclude by the First Derivative Test that the curve has maximal curvature when x = 1/√2.

OpenStudy (camper4834):

rosy don't forget the 7

OpenStudy (camper4834):

7 ln(x)

OpenStudy (camper4834):

:OOOOOO so multiplying your answer by 7 gave me the correct answer that must mean that we can use the curvature of JUST ln(x) as our base

OpenStudy (camper4834):

that doesn't make sense though. doesn't 7 change the function as a whole once we find the curvature formula and its derivative?